# A comprehensive exploration of quasisymmetric stellarators and their coil sets

### v2

@@ 1 Introduction
In recent years, the stellarator community has developed novel numerical optimization algorithms to quickly design stellarators that are quasisymmetric. Quasisymmetry (QS) is a property of the magnetic field, where the field strength on a given magnetic surface can be written as a linear combination of the Boozer angles (Boozer, 1981a) on the surface (Helander, 2014). It can be shown that such fields confine guiding center orbits (Boozer, 1981b; Nührenberg & Zille, 1988; Rodriguez et al., 2020; Burby et al., 2020). Gradient-based optimization algorithms have facilitated the design of magnetic fields with the QS property (Landreman & Paul, 2022), as well as the electromagnetic coils that generate them (Wechsung et al., 2022).
@@ 2 The coil design workflow
To generate multiple devices through the Phases described above, we scan over a number of physics and engineering target characteristics. In terms of physics properties, we target various values of mean rotational transform $\iota$, aspect ratio, and total coil length, while keeping the major radius of 1 meter. The engineering characteristics on the coils impose limits on their curvature and mean squared curvature, which may not exceed $5~{}\text{m}^{-1}$ and $5~{}\text{m}^{-2}$, respectively. The coils are also designed to have uniform incremental arclength. We refer to these as engineering constraints. The goal of these parameter scans is to explore the space of quasisymmetric stellarators that can be generated by coil sets, and observe possible trade-offs between competing physics and engineering characteristics. Whenever ‘quality of quasisymmetry’ for a device is provided in the manuscript, it is calculated by taking the square root of the average of (1) on a number of surfaces of the magnetic field, which we may refer to as $\bar{f}_{\mathrm{QS}}$.
@@ 3 A few QH designs in QUASR
In this section, we present a few devices in the data set, summarized in Table 1 and plotted in Figure 2. We estimate the collisionless guiding-center (Blank, 2004) losses of alpha particles in vacuum magnetic generated the coils of these devices after scaling them up to the ARIES-CS (Najmabadi et al., 2008) minor radius and spawning 5,000 alpha particles isotropically at half-radius in SIMSOPT (Landreman et al., 2021), then tracing until $0.2$ seconds. All devices have favorable particle loss estimates despite deviations from QS, most likely owing to the narrow orbit widths in QH devices (Landreman & Paul, 2022; Buller et al., 2024; Paul et al., 2022). Except for the two field period device, all presented devices have reasonable elongation of flux surfaces in the $RZ$ plane. Note that this measure of shaping (especially when large) can be somewhat misleading in devices with a large helical excursion of the magnetic axis (see Section 5.2 for additional discussion on this). In such cases it is more instructive to measure the elongation in the plane perpendicular to the magnetic axis; for the $n_{\text{fp}}=2$ device, the maximum elongation in such plane reduces to $6.85$, which is still significant. [^1] This quantity was numerically computed by fitting a near-axis model to the field (Section A), and providing the parameters to pyQSC (Landreman et al., 2024).
The first device is a two field period QH field with a ‘crossing point’ with little space for coils, for which coils may nevertheless be found. Although it has the highest particle losses compared to the other devices presented here, the device is reminiscent of the original figure-eight stellarator design (Spitzer, 1958). The $n_{\text{fp}}=4$ device is quite similar to the Landreman-Paul precise QH configuration (Landreman & Paul, 2022), but was discovered here independently with coils. The $n_{\mathrm{fp}}=6$ device shows an example of a configuration with a high number of field periods, which are configurations that have proved challenging to find and provide a set of coils for (Landreman, 2022).
A note must be made on the coils presented for these configurations here. The filamentary coils generally appear quite long, which do not make the coil sets particularly practical. However, having such coils remains useful at least as an initial starting point for further refinement.
@@
ID $n_{\text{fp}}$ mean $\iota$ aspect ratio max. elongation particle losses (%) quality of QS
2229316 2 0.9 6 8.75 3.26 4.03e-04
1655332 3 1.2 6 3.66 0.00 4.22e-04
1630198 4 1.2 8 3.62 0.00 7.54e-04
1641080 5 1.5 8 2.16 0.00 2.26e-03
2008429 6 2.5 12 2.31 0.14 8.09e-04
Table 1: Characteristics of some QH devices from the QUASR data set. The maximum elongation is computed in the $RZ$ plane.
@@ 4 The near-axis quasisymmetry landscape
To construct such a space, we start by providing the elements necessary to characterize a near-axis quasi-symmetric vacuum field (Landreman & Sengupta, 2019; Rodríguez et al., 2023). Such a field is defined by providing the geometry of the magnetic axis (customarily given as a set of Fourier harmonics $\mathbf{R},\mathbf{Z}$ describing the axis shape in cylindrical coordinates), and two scalars $\overline{\eta}$ and $B_{2c}$. Given such inputs, one can evaluate a function $\Delta B_{20}(\mathbf{R},\mathbf{Z},\overline{\eta},B_{2c})$ which measures the magnitude of the symmetry breaking of $B$ at second order in the distance from the magnetic axis (Landreman & Sengupta, 2019; Rodríguez et al., 2023). That is, if $\Delta B_{20}=0$, then the near-axis vacuum field considered has ‘perfect’ quasisymmetry to second order. This scalar function of $(\mathbf{R},\mathbf{Z},\overline{\eta},B_{2c})\in\mathbb{R}^{2N+1}$ variables is too high-dimensional for straightforward plotting. Using the approach in Rodríguez et al. (2023), one may nevertheless reduce this complexity to a minimum that retains the critical structure of the problem. This reduction of the space draws from the following logic:
@@
    the main structural element in the field description is the axis shape, which will remain our free parameters. For visualization purposes, the axis geometry is restricted to a few harmonics (this is key only to visualisation, but not to the concept itself), in this case $\mathbf{R}=(1,R_{n_{\text{fp}}},R_{2n_{\text{fp}}},R_{3n_{\text{fp}}})$, and $\mathbf{Z}=(Z_{n_{\text{fp}}},Z_{2n_{\text{fp}}},Z_{3n_{\text{fp}}})$,
@@
QUASR on the near-axis quasisymmetry phase space
@@
Second, for illustration purposes, we focused on configurations whose axes had 3 dominant harmonics. A device is deemed to have $N$ dominant Fourier harmonics if $R_{k\,n_{\text{fp}}}(1+k^{2}n_{\text{fp}}^{2}),Z_{k\,n_{\text{fp}}}(1+k^{2}n_{\text{fp}}^{2})<0.1$ for $k\geq N+1$. These conditions rely on rescaled Fourier harmonics that are dimensionless and naturally measure relevance of axis harmonics, introduced in Rodriguez et al. (2022b). Around 50% of devices have three dominant harmonics, and this number increases to around 70% when four dominant harmonics are allowed. This hints that the choice to represent magnetic axes in cylindrical coordinates and Fourier harmonics might not be the optimal choice, and an alternative parametrization of the phase space could be better suited (even within the near-axis framework).
@@ 5 Principal component analysis
However, PCA can also have disadvantages. By reducing the dimensionality of the data, we may lose some important information. The transformation to low-dimensional space may distort the geometric relationships that existed in the original space. In addition, PCA is useful when the data is well approximated on a linear manifold of dimension less than $D$, which is certainly not always the case. For example, consider a data set that lies on a three-dimensional helix. The intrinsic dimension of the data is one, since the helix can be unraveled to parameterize any point on it by a single coordinate. But applying PCA to the data set distorts it, since its topology is ill-approximated on a linear manifold of dimension less than three. In cases like this, nonlinear dimensionality reduction techniques can be useful, and the literature on these approaches is rich. For example, isomap (Tenenbaum et al., 2000), and t-Distributed Stochastic Neighbor Embedding (Van der Maaten & Hinton, 2008) are all well-known algorithms to accomplish this task. The disadvantage of nonlinear techniques is that they may have user-defined tuning parameters, which can be difficult to select. Since the underlying topology of our data set is not known, and to avoid the difficulties associated with nonlinear techniques, we use PCA in this work.
@@
    KNN fraction: the fraction of $k$-nearest neighbors in the high-dimensional data set that remain in the two-dimensional projected data points (Lee & Verleysen, 2009; Kobak & Berens, 2019), called the ‘KNN fraction’. The purpose of this quantity is to measure how accurately local relationships are preserved after dimensionality reduction. This is a local measure of accuracy.
@@ 5.1 Feature vectors
$v(\theta)$ $=(1,\cos(2\pi\varphi),\sin(2\pi\varphi),\ldots,\cos(2\pi m_{\text{pol}}\varphi),\sin(2\pi m_{\text{pol}}\varphi)).$
@@
The feature vector associated to a device contains all the Fourier harmonics $\hat{x}_{i,j},\hat{y}_{i,j}$ and $\hat{z}_{i,j}$ associated to the surface, without any additional normalization. [^2] We emphasise this lack of normalization because in the context of PCA it is usually recommended to normalize the data across samples. One common normalization is to rescale each component of the feature vector to a Gaussian with mean zero and covariance one to ensure no feature dominates the analysis due to it being much larger. However, this kind of normalization only makes sense if the different components of the feature vector are not meaningfully related to each other and their values have different orders of magnitude because they come from inherently different quantities or phenomena. In our case, the feature vectors contain Fourier coefficients, which decay as the harmonic increases. Therefore, we chose to forgo the normalization of different Fourier components across the devices because the values of the Fourier coefficients have a very important meaning: they are related to each other to describe a single surface. For the vacuum field devices in QUASR [^3] Additional parameters would be required in the feature vector for devices with plasma pressure., the magnetic field of the device is fully specified by the geometry of a single magnetic surface, regardless of how this is parametrised. This leaves some non-uniqueness in our chosen representation, as different angle choices for the defining surface can lead to different fourier coefficients defining the same magnetic surface geometry. To remove this indeterminacy, we choose the angles on the surface $(\varphi,\theta)$ to correspond to Boozer coordinates. Even then, there remains certain non-uniqueness; for example, the same configuration rotated by a half-period about the $Z$-axis, or reflected about the $XY$ plane would correspond to different feture vectors. Thus, to ensure that geometrically similar devices are close to each other in Fourier space, in addition to
@@
A more compact feature vector could have been generated by converting the representation in (5)-(9) to a cylindrical one, thereby reducing the number of Fourier series from three to two. We do not do so here because the coil design workflow relies on the Cartesian representation of the surfaces. [^4] Although Cartesians are used in the heart of the coil design workflow, this does not mean any field shape is possible. Note that in Phase I (the near-axis phase of the coil design) it is assumed that the magnetic axis can be represented in cylindrical coordinates. As a result, we will not find devices with more general magnetic axis geometries, e.g., devices where the axis is anywhere vertical, or knotted. Generalizing our implementation is the subject of future work. In addition, this conversion is not lossless when restricted to a finite dimensional Fourier series, as explained in section 5.2. Other forms of parametrisation could also be possible. It is tempting, for example, to represent each configuration by its second order near-axis model, which constitutes a significantly reduced space. Although this could be interesting, it does not correspond to a one-to-one map, and thus should not be used as a unique representation. [^5] The main reason behind this is that, a priori, one could devise an infinity of distinct fields, all of which correspond to the same second order near-axis construction. Another alternative could be to use a vector of Fourier harmonics associated to the coils and their currents. We did not do this due to the inherent non-uniqueness of coil sets, i.e., multiple, vastly different coil sets can produce similar magnetic fields. That being said, we have not fully explored these directions, and it might be interesting to do this in the future.
@@
Each point on the landscape plot corresponds to a specific vacuum magnetic field, with associated magnetic surfaces. There are box-shaped regions delineated by dashed lines on the plot. Taking the point at the center of each box, we compute cross sections of the associated magnetic surface. The cross sections are plotted in black, red, and green, associated to the cylindrical angle $\phi=0,\pi/(2n_{\text{fp}}),\pi/n_{\text{fp}}$, respectively. This figure makes it clear how devices differ from each other on the plane, and importantly how moving along the principal components affects the geometry of the device. Moving in the positive direction associated to $\mathbf{d}^{*}_{1}$, the elongation appears to increase, while going in the negative direction the elongation decreases. In the positive $\mathbf{d}^{*}_{2}$ direction, the relative positions of the cross sections move, and in particular the axis excursion increases. These cross sections do not necessarily correspond to physically relevant devices, as illustrated by the leftmost column, where the cross sections are self-intersecting. This is not surprising, as the devices lie on a nonlinear manifold, and the PCA slices through it with a linear one. If one strays too far away from the origin of the plot (the sample mean), the approximation errors become noticeable and unphysical devices appear.
Within the space populated by the QUASR devices, we find that converting the highly-shaped surfaces, e.g. the device in Figure 7, to a cylindrical representation, as is used in the more standard equilibrium solvers VMEC or DESC, requires more than $m_{\text{pol}},n_{\text{tor}}=20$ Fourier modes. This is significantly larger than the number of modes used to design the stellarator ($m_{\text{pol}},n_{\text{tor}}=10$) with the representation in equation (5)-(9), which could difficult describing these configurations by the more standard means. Figure 7 illustrates the geometrical origin of this difficulty, where due to the large helical excursion of the magnetic axis, cross sections in the $RZ$ plane become highly elongated. This occurrence is particularly true for the devices on the right branch, in the bottom quadrant, where standard equilibrium solvers such as VMEC experience difficulties. [^6] These problems are probably shared by DESC (Panici et al., 2023) which uses the same cylindrical representation, but could perhaps be alleviated in the Frenet version of the GVEC code (Hindenlang et al., 2024). This can also lead to a misleading interpretation of the elongation of the device.
@@
Figure 11: The nonlinear manifold in Figure 10 can be parametrized with a single coordinate using the isomap algorithm. The row of devices are uniformly sampled from the isomap embedding, and the numbers below these devices corresponds to the isomap coordinate and the quality of quasisymmetry.

### v3

@@ 1 Introduction
In recent years, the stellarator community has developed novel numerical optimization algorithms to quickly design stellarators that are quasisymmetric. Quasisymmetry (QS) is a property of the magnetic field, where the field strength on a given magnetic surface can be written as a linear combination of the Boozer angles (Boozer, 1981a) on the surface (Helander, 2014). It can be shown that such fields confine guiding center orbits (Boozer, 1981b; Nührenberg & Zille, 1988; Rodriguez et al., 2020; Burby et al., 2020). Gradient-based optimization algorithms have facilitated the design of magnetic fields with the QS property (Landreman & Paul, 2022), as well as the electromagnetic coils that generate them (Wechsung et al., 2022b).
@@ 2 The coil design workflow
To generate multiple devices through the phases described above, we scan over a number of physics and engineering target characteristics. In terms of physics properties, we target various values of mean rotational transform $\iota$, aspect ratio, and total coil length, while keeping the major radius of 1 meter. The engineering characteristics on the coils impose limits on their curvature and mean squared curvature, which may not exceed $5~{}\text{m}^{-1}$ and $5~{}\text{m}^{-2}$, respectively. The coils are also designed to have uniform incremental arclength. We refer to these as engineering constraints. The goal of these parameter scans is to explore the space of quasisymmetric stellarators that can be generated by coil sets, and observe possible trade-offs between competing physics and engineering characteristics. Whenever ‘quality of quasisymmetry’ for a device is provided in the manuscript, it is calculated by taking the square root of the average of (1) on a number of surfaces of the magnetic field, which we may refer to as $\bar{f}_{\mathrm{QS}}$. Note that by construction, the surfaces on which we optimize for quasisymmetry in the final stage of the workflow are tangent the magnetic field at a fixed number of quadrature points on the surface. As a result, $\mathbf{B}\cdot\mathbf{n}$ error is numerically zero when queried at those same quadrature points. This is notably different from the approach in (Jorge et al., 2023), where the $\mathbf{B}\cdot\mathbf{n}$ error is explicitly minimized.
The stellarator design problem amounts to a constrained optimization problem, where the ODE (phase 1) or PDE (phase 2, 3) constraint is satisfied to numerical precision at each iteration of the optimization loop. This was not necessarily the case for the physics and engineering constraints. However, all the target physics and engineering constraints were satisfied to a precision of $0.1\%$ by using a penalty method. After a fixed budget of iterations of the minimization algorithm, we increased by a factor of 10 the penalty weights associated to the constraints that were violated by more than this accuracy threshold. Then, with these new set of weights, the minimization algorithm is restarted. This ensures that these constraints are satisfied to the desired accuracy at the end of the optimization workflow. After a fixed number of restarts, devices that do not satisfy the engineering and physics constraints to $0.1\%$ accuracy, even after the penalty reweighting, are discarded.
@@ 3 A few QH designs in QUASR
In this section, we present a few devices in the data set, summarized in Table 1 and plotted in Figure 2. The table also includes a link to a device manifest page, where additional physics and engineering information is provided. We estimate the collisionless guiding-center (Blank, 2004) losses of alpha particles in vacuum magnetic generated the coils of these devices after scaling them up to the ARIES-CS (Najmabadi et al., 2008) minor radius and spawning 5,000 alpha particles isotropically at half-radius in SIMSOPT (Landreman et al., 2021), then tracing until $0.2$ seconds. All devices have favorable particle loss estimates despite deviations from QS, most likely owing to the narrow orbit widths in QH devices (Landreman & Paul, 2022; Buller et al., 2024; Paul et al., 2022). Except for the two field period device, all presented devices have reasonable elongation of flux surfaces in the $RZ$ plane. Note that this measure of shaping (especially when large) can be somewhat misleading in devices with a large helical excursion of the magnetic axis (see Section 5.2 for additional discussion on this). In such cases it is more instructive to measure the elongation in the plane perpendicular to the magnetic axis; for the $n_{\text{fp}}=2$ device, the maximum elongation in such plane reduces to $6.85$, which is still significant.This quantity was numerically computed by fitting a near-axis model to the field (Section A), and providing the parameters to pyQSC (Landreman et al., 2024).
The first device is a two field period QH field with a ‘crossing point’ with little space for coils, for which coils may nevertheless be found. Although it has the highest particle losses compared to the other devices presented here, the device is reminiscent of the original figure-eight stellarator design (Spitzer, 1958).
[…]
The $n_{\mathrm{fp}}=6$ device shows an example of a configuration with a high number of field periods, which are configurations that have proved challenging to find and provide a set of coils for (Landreman, 2022).
A note must be made on the coils presented for these configurations here. The filamentary coils sometimes appear quite long, which might not make the coil sets particularly practical. However, having such coils remains useful at least as an initial starting point for further refinement.
@@
ID $n_{\text{fp}}$ mean $\iota$ aspect ratio max. elongation losses (%) quality of QS device manifest
2229316 2 0.9 6 8.75 3.26 4.03e-04 [link]
1655332 3 1.2 6 3.66 0.00 4.22e-04 [link]
1630198 4 1.2 8 3.62 0.00 7.54e-04 [link]
1641080 5 1.5 8 2.16 0.00 2.26e-03 [link]
2008429 6 2.5 12 2.31 0.14 8.09e-04 [link]
Table 1: Characteristics of some QH devices from the QUASR data set. The maximum elongation is computed in the $RZ$ plane. The final column of the table contains links to a device manifest page that contain more physics and engineering information about each design.
@@ 4 The near-axis quasisymmetry landscape
To construct such a space, we start by providing the elements necessary to characterize a near-axis quasi-symmetric vacuum field (Landreman & Sengupta, 2019; Rodríguez et al., 2023). Such a field is defined by providing the geometry of the magnetic axis (customarily given as a set of Fourier harmonics $\mathbf{R},\mathbf{Z}$ describing the axis shape in cylindrical coordinates), and two scalars $\overline{\eta}$ and $B_{2c}$. Given such inputs, one can evaluate a function $\Delta B_{20}(\mathbf{R},\mathbf{Z},\overline{\eta},B_{2c})$ which measures the magnitude of the symmetry breaking of $B$ at second order in the distance from the magnetic axis (Landreman & Sengupta, 2019; Rodríguez et al., 2023). That is, if $\Delta B_{20}=0$, then the near-axis vacuum field considered has ‘perfect’ quasisymmetry to second order. This scalar function of $(\mathbf{R},\mathbf{Z},\overline{\eta},B_{2c})\in\mathbb{R}^{2N+1}$ variables is too high-dimensional for straightforward plotting. Using the approach in Rodríguez et al. (2023), one may nevertheless reduce the dimensionality in such a way that the critical structure of the problem is retained. This reduction of the space draws from the following logic:
@@
    the main structural element in the field description is the axis shape, which will remain our free parameters. For visualization purposes, the axis geometry is restricted to a few harmonics (this is key only to visualization, but not to the concept itself), in this case $\mathbf{R}=(1,R_{n_{\text{fp}}},R_{2n_{\text{fp}}},R_{3n_{\text{fp}}})$, and $\mathbf{Z}=(Z_{n_{\text{fp}}},Z_{2n_{\text{fp}}},Z_{3n_{\text{fp}}})$,
@@
QUASR on the near-axis quasisymmetry landscape
@@
Second, for illustration purposes, we focused on configurations whose axes had 3 dominant harmonics. A device is deemed to have $N$ dominant Fourier harmonics if $R_{k\,n_{\text{fp}}}(1+k^{2}n_{\text{fp}}^{2}),Z_{k\,n_{\text{fp}}}(1+k^{2}n_{\text{fp}}^{2})<0.1$ for $k\geq N+1$. These conditions rely on rescaled Fourier harmonics that are dimensionless and naturally measure relevance of axis harmonics, introduced in Rodriguez et al. (2022b). Around 50% of devices have three dominant harmonics, and this number increases to around 70% when four dominant harmonics are allowed. During the globalization of the near-axis phase of the workflow, the axis is described only using two Fourier harmonics. Then, the solution is subsequently polished and the number of harmonics used to describe the axis is increased. This continuation procedure may have influenced these statistics. Additionally, note that the choice to represent magnetic axes in cylindrical coordinates and Fourier harmonics might not be the optimal choice, and an alternative parametrization of the phase space could be better suited (even within the near-axis framework).
@@ 5 Principal component analysis
However, PCA can also have disadvantages. By reducing the dimensionality of the data, we may lose some important information. The transformation to low-dimensional space may distort the geometric relationships that existed in the original space. In addition, PCA is useful when the data is well approximated on a linear manifold of dimension less than $D$, which is certainly not always the case. For example, consider a data set that lies on a three-dimensional helix. The intrinsic dimension of the data is one, since the helix can be unraveled to parameterize any point on it by a single coordinate. But applying PCA to the data set distorts it, since its topology is ill-approximated on a linear manifold of dimension less than three. In cases like this, nonlinear dimensionality reduction techniques can be useful, and the literature on these approaches is rich. For example, isomap (Tenenbaum et al., 2000), and t-Distributed Stochastic Neighbor Embedding (Van der Maaten & Hinton, 2008) are all well-known algorithms to accomplish this task. The disadvantage of nonlinear techniques is that they may have user-defined tuning parameters, which can be difficult to select. Since the underlying topology of our data set is not known, and to avoid the difficulties associated with nonlinear techniques, we use PCA in this work. Auto encoders have also been proposed as interesting methods for compression or dimensionality reduction. (Bourlard & Kamp, 1988) showed that auto encoders perform a similar operation to SVD or low rank matrix approximation. Since PCA is a type of low rank matrix approximation, this is an interesting connection to what is done in this work.
@@
    KNN fraction: the fraction of $k$-nearest neighbors in the high-dimensional data set that remain in the two-dimensional projected data points (Lee & Verleysen, 2009; Kobak & Berens, 2019), called the ‘KNN fraction’. The purpose of this quantity is to measure how accurately local relationships are preserved after dimensionality reduction. Note that for very large $k$, this fraction approaches 1, so some reasonable choice of $k$ has to be picked. This is a local measure of accuracy.
@@ 5.1 Feature vectors
$v(\theta)$ $=(1,\cos(2\pi\theta),\sin(2\pi\theta),\ldots,\cos(2\pi m_{\text{pol}}\theta),\sin(2\pi m_{\text{pol}}\theta)).$
@@
The feature vector associated to a device contains all the Fourier harmonics $\hat{x}_{i,j},\hat{y}_{i,j}$ and $\hat{z}_{i,j}$ associated to the surface, without any additional normalization. We emphasize this lack of normalization because in the context of PCA it is usually recommended to normalize the data across samples. One common normalization is to rescale each component of the feature vector to a Gaussian with mean zero and covariance one to ensure no feature dominates the analysis due to it being much larger. However, this kind of normalization only makes sense if the different components of the feature vector are not meaningfully related to each other and their values have different orders of magnitude because they come from inherently different quantities or phenomena. In our case, the feature vectors contain Fourier coefficients, which decay as the harmonic increases. Therefore, we chose to forgo the normalization of different Fourier components across the devices because the values of the Fourier coefficients have a very important meaning: they are related to each other to describe a single surface. For the vacuum field devices in QUASR, the magnetic field of the device is fully specified by the geometry of a single magnetic surface, regardless of how this is parametrised. Additional parameters would be required in the feature vector for devices with plasma pressure, and normalization might be indicated there. This leaves some non-uniqueness in our chosen representation, as different angle choices for the defining surface can lead to different fourier coefficients defining the same magnetic surface geometry. To remove this indeterminacy, we choose the angles on the surface $(\varphi,\theta)$ to correspond to Boozer coordinates. Even then, there remains certain non-uniqueness; for example, the same configuration rotated by a half-period about the $Z$-axis, or reflected about the $XY$ plane would correspond to different feture vectors. Thus, to ensure that geometrically similar devices are close to each other in Fourier space, in addition to
@@
A more compact feature vector could have been generated by converting the representation in (5)-(9) to a cylindrical one, thereby reducing the number of Fourier series from three to two. We do not do so here because the coil design workflow relies on the Cartesian representation of the surfaces. Although Cartesian coordinates are used in the coil design workflow, this does not mean any field shape is possible. Note that in Phase I (the near-axis phase of the coil design) it is assumed that the magnetic axis can be represented in cylindrical coordinates. As a result, we will not find devices with more general magnetic axis geometries, e.g., devices where the axis is anywhere vertical, or knotted. Generalizing our implementation is the subject of future work.
In addition, this conversion is not lossless when restricted to a finite dimensional Fourier series, as explained in section 5.2. Other forms of parametrization could also be possible. It is tempting, for example, to represent each configuration by its second order near-axis model, which constitutes a significantly reduced space. Although this could be interesting, it does not correspond to a one-to-one map, and thus should not be used as a unique representation. This is because, a priori, one could devise an infinity of distinct fields, all of which correspond to the same second order near-axis construction. Another alternative could be to use a vector of Fourier harmonics associated to the coils and their currents. We did not do this due to the inherent non-uniqueness of coil sets, i.e., multiple, vastly different coil sets can produce similar magnetic fields. That being said, we have not fully explored these directions, and it might be interesting to do this in the future.
@@
Each point on the landscape plot corresponds to a specific vacuum magnetic field, with associated magnetic surfaces. There are box-shaped regions delineated by dashed lines on the plot. Taking the point at the center of each box, we compute cross sections of the associated magnetic surface. The cross sections are plotted in black, red, and green, associated to the cylindrical angle $\phi=0,\pi/(4n_{\text{fp}}),\pi/(2n_{\text{fp}})$, respectively. This figure makes it clear how devices differ from each other on the plane, and importantly how moving along the principal components affects the geometry of the device. Moving in the positive direction associated to $\mathbf{d}^{*}_{1}$, the elongation appears to increase, while going in the negative direction the elongation decreases. In the positive $\mathbf{d}^{*}_{2}$ direction, the relative positions of the cross sections move, and in particular the axis excursion increases. These cross sections do not necessarily correspond to physically relevant devices, as illustrated by the leftmost column, where the cross sections are self-intersecting. This is not surprising, as the devices lie on a nonlinear manifold, and the PCA slices through it with a linear one. If one strays too far away from the origin of the plot (the sample mean), the approximation errors become noticeable and unphysical devices appear.
Within the space populated by the QUASR devices, we find that converting the highly-shaped surfaces, e.g. the device in Figure 7, to a cylindrical representation, as is used in the more standard equilibrium solvers VMEC or DESC, requires more than $m_{\text{pol}},n_{\text{tor}}=20$ Fourier modes. This is significantly larger than the number of modes used to design the stellarator ($m_{\text{pol}},n_{\text{tor}}=10$) with the representation in equation (5)-(9). Figure 7 illustrates the geometrical origin of this difficulty, where due to the large helical excursion of the magnetic axis, cross sections in the $RZ$ plane become highly elongated, leading to a misleading interpretation of the elongation of the device. This occurrence is particularly true for the devices on the right branch, in the bottom quadrant, where standard equilibrium solvers such as VMEC experience difficulties. These problems are probably shared by DESC (Panici et al., 2023) which uses the same cylindrical representation, but could perhaps be alleviated in the Frenet version of the GVEC code (Hindenlang et al., 2024).
@@
Figure 11: The nonlinear manifold in Figure 10 can be parametrized with a single coordinate using the isomap algorithm. The row of devices are uniformly sampled from the isomap embedding, and the numbers below these devices corresponds to the isomap coordinate and the quality of quasisymmetry. The color on the magnetic surfaces corresponds to the local field strength